# A Universal Framework for Horizon-Scale Tests of Gravity with Black Hole Shadows

### v1

@@ Abstract
In this Letter, we have developed a numerically efficient framework for evaluating parameters in metric theories of gravity, and applied it to constrain the horizon-scale magnetic field in the Kerr–Bertotti–Robinson (Kerr-BR) spacetime using the latest EHT observations. The method’s adaptive ray-tracing strategy achieves near-linear computational efficiency without loss of numerical accuracy. Owing to this efficiency, the framework enables high precision shadow modeling at minimal computational cost and, for the first time, supports statistically robust inference of black hole parameters from horizon-scale observations for arbitrary stationary black holes. The above framework is applied to the recently obtained Kerr–BR black hole, an exact magnetized and rotating solution to the Einstein field equations. We have evaluated the horizon-scale magnetic fields of M87* and Sgr A*, with the latter showing a field strength of $93.3^{+14.7}_{-23.8}\mathrm{G}$, consistent with the equipartition estimate of $71\mathrm{G}$ from polarized ALMA observations, thereby supporting Einstein’s gravity.
@@
Introduction.—In a major step forward, the Event Horizon Telescope (EHT) Collaboration has released last month polarized images of M87* at 230 GHz, obtained during the 2017, 2018, and 2021 observing campaigns Akiyama et al. (2025). The inclusion of the 12-m Kitt Peak Telescope and the Northern Extended Millimetre Array in 2021 substantially improved the baseline coverage of the EHT, yielding a more precise measurement of the M87* ring diameter of $43.9\pm 0.6~\mu{\rm as}$, compared to the $42\pm 3~\mu{\rm as}$ value derived from the 2017 observations and published by the EHT Collaboration in 2019 Akiyama et al. (2019a, b, c, d). These unprecedented observations have, in turn, motivated extensive theoretical and numerical efforts to interpret the ring morphology and polarization signatures Psaltis et al. (2020); Cunha and Herdeiro (2020); Crinquand et al. (2022); Davelaar and Haiman (2022); Cunha et al. (2023); Galishnikova et al. (2023); Chen et al. (2025); Uniyal et al. (2025). In particular, the remarkable contributions from several groups Cunha and Herdeiro (2018); Mizuno et al. (2018); Gralla et al. (2019); Perlick and Tsupko (2022); Bronzwaer and Falcke (2021); Konoplya and Zhidenko (2021); Vagnozzi et al. (2023); Chen et al. (2023); Ayzenberg et al. (2025), most notably the development of ray-tracing techniques Cunha et al. (2015, 2019); Hu et al. (2021); Bacchini et al. (2021); Zhang et al. (2025), have enabled the imaging of non-integrable and fully numerical black-hole spacetimes, providing a crucial bridge between theory and EHT observations.
@@
However, high resolution backward ray-tracing simulations are notoriously computationally intensive, making it difficult to apply statistical inference methods such as Markov Chain Monte Carlo (MCMC) directly to constrain model parameters from observational data Johannsen et al. (2016). To overcome this limitation, we construct a general framework that unifies numerical imaging and parameter estimation in metric theories of gravity by developing an adaptive backward ray-tracing algorithm inspired by the adaptive mesh refinement techniques in numerical relativity Clough et al. (2015); Buschmann et al. (2022), which reduces the computational scaling from geometric to nearly linear, and combining it with a Schwarzschild-calibrated normalization scheme that maps simulated image sizes directly to physical observables, independently of the observer’s configuration. This enables high precision shadow modeling at minimal computational cost and, for the first time, allows statistically robust inference of black hole parameters from horizon-scale observations for arbitrary black hole spacetimes.
@@
In this Letter, we apply the above framework to the recently obtained Kerr–Bertotti–Robinson (Kerr–BR) black hole, an exact magnetized and rotating solution to the Einstein field equations Podolsky and Ovcharenko (2025). Unlike the Kerr-Melvin spacetimes Ernst and Wild (1976), the Kerr–BR spacetime belongs to the Petrov type-D class and features finite, asymptotically uniform electromagnetic fields and bounded ergoregions, thereby providing a more realistic model for astrophysical black holes embedded in external magnetic environments Ovcharenko and Podolský (2025); Wang (2025); Wang et al. (2025); Ali and Ghosh (2025); Zeng and Wang (2025); Zhang and Wei (2025). By matching the theoretical predictions of the Kerr–BR spacetime with the latest EHT observations of M87* and Sgr A*, quantitative constraints are placed on the magnetic field strength. In particular, the analysis yields a magnetic field strength of $93.3^{+14.7}_{-23.8}\mathrm{G}$ for Sgr A*, consistent with the equipartition value $71\mathrm{G}$ inferred from polarized ALMA observations of submillimeter flares Michail et al. (2023). Our results suggest that the magnetized environment can be consistently described by the exact Kerr–BR solution within GR, without invoking any modification to the theory.
@@
with $I_{1}=1-B^{2}a^{2}/2$ and $I_{2}=1-B^{2}a^{2}$. The metric reduces to the Kerr solution in the limit $B=0$, but not to the Schwarzschild spacetime for $a=0$, reflecting the intrinsic coupling between rotation and the external magnetic field. Here $M$ is an integration constant interpreted as the mass parameter of the black hole. The outer and inner horizons of the black hole are determined by the roots of the metric function $\Delta\!=\!0$, namely $r_{\pm}\!=\!\tfrac{MI_{1}I_{2}\pm I_{1}\sqrt{M^{2}I_{2}-a^{2}I_{1}^{2}}}{I_{1}^{2}-B^{2}M^{2}I_{2}}$. The explicit form of the electromagnetic field is given in Ref. Podolsky and Ovcharenko (2025), but it is not essential to our present analysis; we are concerned only with its overall strength. In this Letter, our goal is to estimate the magnetic field strength of realistic supermassive black holes by developing a new inference framework based on adaptive ray-tracing simulations, which directly link the observed shadow morphology to the underlying uniform magnetic field.
Photon trajectories are integrated using the Hamiltonian formulation of null geodesics in the Kerr–BR spacetime,
@@
where $p_{\mu}=g_{\mu\nu}p^{\nu}$ and $p^{\mu}=(\dot{t},\dot{r},\dot{\vartheta},\dot{\varphi})$. The Killing vectors $\xi^{\mu}$ and $\psi^{\mu}$ yield two conserved quantities, the energy $\mathcal{E}$ and axial angular momentum $L_{z}$ Wald (1984), which allow the null geodesic equations to be numerically integrated in our backward ray-tracing scheme. Further computational details are given in Refs. Liu et al. (2025a, 2024a, b, 2024b, c). As illustrated in Fig. 1, a virtual observer is placed at a large radial distance, where the spacetime is effectively flat. The observer’s image plane, spanned by Cartesian coordinates $(x,y)$, is divided into an adaptive grid representing different initial photon directions within a field of view $\alpha_{\rm fov}$. Each grid point $(m,n)$ corresponds to a null geodesic that is integrated backward in time until the photon either crosses the event horizon or escapes to infinity. In our simulations, these two outcomes are distinguished on the image plane: escaping photons are marked by red pixels, while captured photons are marked by purple pixels. After all grid points have been traced, the collection of purple pixels forms the simulated black hole shadow, whose outer boundary is illustrated by the purple ring in Fig. 1. For quantitative analysis, we introduce a reference circle, shown as the orange ring in Fig. 1, to compare with the noncircular shadow. The circle is defined by the circumcircle passing through the characteristic points $P_{A}(x_{A},y_{A})$ [^1] The normalized coordinates $(x_{i},y_{i})\in(0,1]$ are defined by the image resolution, e.g., $x_{A}=m/n$ for the $m$-th pixel in an $n\times n$ grid., $P_{B}(x_{B},y_{B})$, and $P_{C}(x_{C},y_{C})$, while $P_{D}(x_{D},y_{D})$ serves as a key indicator of the shadow’s deformation. Let the coordinates of the circumcenter be denoted by $P_{O}(x_{O},y_{O})$. The center satisfies the following system of equations:
@@
Through numerical backward ray-tracing simulations in the Kerr–BR spacetime, we establish explicit relationships between the observable quantities, the black hole parameters, and the observer’s configuration. Accordingly, one can express $R_{s}\!\equiv\!R_{s}(\psi)$ and $\mathcal{D}_{s}\!\equiv\!\mathcal{D}_{s}(\psi)$ with $\psi\!\equiv\!\{\tilde{a},\tilde{B},\vartheta_{0}\}$, where $\tilde{a}\!=\!a/M$ and $\tilde{B}\!=\!BM$ denote the dimensionless spin and magnetic field strength, and $\vartheta_{0}$ is the inclination angle. For each parameter set $(\tilde{a},\tilde{B},\vartheta_{0})$, we ray-trace a two-dimensional shadow image on the observer’s image plane and extract the image-plane observables $(R_{s},\mathcal{D}_{s})$ from its pixels. Because $R_{s}$ is an image-plane quantity, we convert it into an apparent angular radius and apply a Schwarzschild-calibrated, so that $\theta_{s}\propto R_{s}$; which details are provided in Appendix A. With this calibration, the mappings $\theta_{s}(\psi)$ and $\mathcal{D}_{s}(\psi)$ provide a direct link between $\psi$ and the observable shadow features. Using EHT measurements, we can constrain the black hole parameters $\tilde{a}$, $\tilde{B}$ and $\vartheta_{0}$. To achieve this, we construct the likelihood function as
@@
when both observed $\theta_{s}$ and $\mathcal{D}_{s}$ are available. Here, $D_{\mathrm{tot}}\!\equiv\!\{\theta_{\rm EHT},\sigma_{\theta},\mathcal{D}_{\rm EHT},\sigma_{\mathcal{D}}\}$ denotes a set of EHT observational data. In general, the probability densities $\mathcal{P}_{\theta_{s}}$ and $\mathcal{P}_{\mathcal{D}_{s}}$ are assumed to follow Gaussian distributions centered at the observed mean values $\theta_{\rm EHT}$ and $\mathcal{D}_{\rm EHT}$ with standard deviations $\sigma_{\theta}$ and $\sigma_{\mathcal{D}}$, respectively, when the reported uncertainties are symmetric, whereas they are modeled as piecewise Gaussians with different standard deviations on either side of the mean value when the uncertainties are asymmetric. Then, the posterior distribution of theoretical parameters is obtained via Bayes’ Theorem: $\mathcal{P}(\psi|D_{\mathrm{tot}})\propto\mathcal{L}(D_{\mathrm{tot}}|\psi)~\Pi(\psi)$, where $\Pi(\psi)$ is the prior distribution of theoretical parameters, which is taken to be uniform. We employ the MCMC method to sample from this posterior distribution and finally obtain the constraints on parameters $\tilde{a}$, $\tilde{B}$, and $\vartheta_{0}$.
Our overall modeling strategy is conceptually straightforward, but its implementation involves several key technical challenges. The first concerns how to map the pixel coordinates of the characteristic points $P_{A}$, $P_{B}$, $P_{C}$, and $P_{D}$ onto physically meaningful quantities relevant to astronomical observations, while the second relates to maintaining numerical precision and computational efficiency. Since the accuracy of the extracted shadow features is highly sensitive to the image resolution, as illustrated in Fig. 2 for the Kerr case where the axial ratio $\mathcal{D}_{s}$ varies with spin, showing that numerical convergence is achieved only at sufficiently high resolutions and the curves become progressively smoother. For this, we developed an adaptive backward ray-tracing algorithm that performs high-resolution calculations only in the vicinity of the critical pixels. This adaptive scheme reduces the computational cost from exponential to nearly linear scaling with resolution, enabling us to explore the black hole parameter space at extremely high resolutions. The algorithmic details and validation tests of this method are presented in Appendix A.
@@
Results.—Having established a numerically consistent and efficient framework for shadow modeling, we next apply it to the EHT targets M87 and Sgr A*, linking their observed shadow features to theoretical predictions within the Kerr–BR spacetime and enabling a quantitative inference of the underlying magnetic field strength $\tilde{B}$. To facilitate the comparison between observations and simulations, the relevant observational quantities and corresponding model parameters are listed in Table 1, with the latter expressed in geometrized units (c=G=M=1)(c\!\!=\!\!G\!\!=\!\!M\!\!=\!\!1).
@@
For the two EHT targets, M87* and Sgr A*, substituting the corresponding parameters from Table 1 into the above expression yields the theoretical angular radii $\theta_{s}$, which serve as the model predictions for subsequent parameter inference. Before comparing these theoretical predictions with observations, it is essential to ensure the numerical convergence of the simulated shadow contours. In general, the simulated shadow contours corresponding to different parameter choices converge at comparable resolutions, as illustrated in Fig. 2. Accordingly, in evaluating $\theta_{s}$ and $\mathcal{D}_{s}$, we find that an effective resolution of 4K (i.e., $4096\times 4096$ pixels) is sufficient to ensure full numerical convergence. With the numerical accuracy thus guaranteed, we proceed to constrain the model parameters using EHT observations. To ensure sufficient sampling of the parameter space, we generated $10^{4}$ simulated shadow images for each of M87* and Sgr A* under randomly selected combinations of $(\tilde{a},\tilde{B},\vartheta_{0})$. The resulting dataset was interpolated to obtain continuous mappings of $(\theta_{s},\mathcal{D}_{s})$, which were then used to calculate the posterior distributions of theoretical parameters.
Figure 3: Corner plot showing the the one- and two-dimensional posterior probability distributions of the BH parameters for M87*, based on the three-epoch combined measurements from the EHT Collaboration Akiyama et al. (2025). The dashed lines indicate the $68\%$ CL, while the dot-dashed line marks the mean value of $\tilde{B}$.
@@
In Fig. 3, we show the one- and two-dimensional posterior distributions of parameters for M87*, based on the three-epoch combined measurements of the EHT Collaboration Akiyama et al. (2025). The magnetic field strength is constrained to $BM\!=\!0.79^{+0.12}_{-0.11}\times 10^{-11}$. The posterior peaks at high spin, a/M>0.759a/M\!>\!0.759, consistent with earlier jet-power arguments Akiyama et al. (2019d). The inclination angle remains weakly constrained, with $\vartheta_{0}\!>\!\frac{77\pi}{250}$ at 68% confidence level (CL). Fig. 4 presents the corresponding posterior distributions for Sgr A*, derived using the EHT 2022 observational data Akiyama et al. (2022). In this case, the term $\mathcal{P}_{\mathcal{D}_{s}}$ in equation (6) is omitted, since only $\theta_{s}$ is available (see Table 1). The inferred magnetic field strength is higher, $BM\!=\!1.53^{+0.22}_{-0.17}\times 10^{-11}$, while the correlations among $a/M$, BMBM, and $\vartheta_{0}$ appear marginal. This is not due to larger observational uncertainties but rather because the Sgr A* dataset provides no effective constraint on the shadow Axial ratio $\mathcal{D}_{s}$, leaving one fewer observable to break parameter degeneracies. Nevertheless, both analyses consistently favor the presence of a nonzero uniform magnetic field in the Kerr–BR spacetime, with the M87* results yielding a tighter credible interval for the field strength.
@@
$B_{\rm Gauss}=\frac{c^{4}}{G_{N}^{3/2}M}B=2.36\times 10^{19}\left(\frac{M_{\odot}}{M}\right)(BM)\ {\rm G},$ (7)
where $M$ denotes the black hole mass, and $B_{\rm Gauss}$ is expressed in gauss. From the posterior distributions of $\tilde{B}$ and equation (7), we derive $B_{\rm Gauss}\!=\!0.0294^{+0.0045}_{-0.0055}\,{\rm G}$ (M87*) and $B_{\rm Gauss}\!=\!93.3^{+14.7}_{-23.8}\,{\rm G}$ (Sgr A*), both at 68% CL. Remarkably, the inferred field strength of Sgr A* agrees well with the equipartition magnetic field ($B_{\rm Gauss}\!\sim\!71{\rm G}$) derived from polarized ALMA observations of submillimeter flares Michail et al. (2023), while the much weaker field of M87* reflects the expected $1/M$ scaling of horizon-scale magnetization. This consistency demonstrates that the Kerr-BR solution, an exact magnetized black hole spacetime within GR, when constrained purely by shadow geometry and EHT priors, can naturally reproduce independent polarimetric estimates of magnetic fields in systems differing by over three orders of magnitude in mass. It thus shows that horizo-scale magnetism can be self consistently explained within GR, without invoking any modification of Einstein’s theory, providing a unified geometric route to quantifying magnetic fields around both weakly and strongly magnetized black holes.
@@
Finally, we summarize all inferred constraints in Table 2. In addition to the results presented in Figs. 3 and 4, we incorporate the other EHT measurements of M87* for comparison, including the 2018 data reanalysis published in 2024 Akiyama et al. (2024) and the most recent three-epoch joint analysis combining the 2017, 2018, and 2021 observation Akiyama et al. (2025). The latter study, which added the 12-m Kitt Peak Telescope and the NOEMA array, provides substantially improved baseline coverage and confirms the long-term stability of the horizon-scale ring across all epochs. Taken together, these results demonstrate that Kerr-BR model yields magnetic field strengths consistent with the latest EHT constraints while remaining fully compatible with GR, thereby reinforcing the viability of horizon-scale magnetism as a purely geometric observable.

### v2

@@ Abstract
In this Letter, we develop a numerically efficient framework for evaluating parameters in metric theories of gravity, and apply it to constrain the horizon-scale magnetic field in the Kerr–Bertotti–Robinson spacetime using the latest Event Horizon Telescope observations. The method’s adaptive ray-tracing strategy achieves near-linear computational efficiency without loss of numerical accuracy. This efficiency allows for high-precision black hole shadow modeling at a minimal computational cost and, for the first time, supports statistically robust parameter inference for arbitrary stationary black holes from horizon-scale observations. Applying the framework to the exact magnetized and rotating Kerr–Bertotti–Robinson solution [Phys. Rev. Lett. $\textbf{135},181401(2025)$], we evaluate the horizon-scale magnetic fields of M87* and Sgr A*. The analysis yields a field strength of $93.3^{+14.7}_{-23.8}\,\mathrm{G}$ for Sgr A*, consistent with the $71\,\mathrm{G}$ equipartition estimate from polarized ALMA observations, thereby supporting the predictions of Einstein’s gravity.
@@
Introduction.—In a major step forward, the Event Horizon Telescope (EHT) Collaboration has recently released polarized images of M87* at 230 GHz, obtained during the 2017, 2018, and 2021 observing campaigns Akiyama et al. (2025). The inclusion of the 12-m Kitt Peak Telescope and the Northern Extended Millimetre Array in 2021 substantially improved the baseline coverage of the EHT, yielding a more precise measurement of the M87* ring diameter of $43.9\pm 0.6~\mu{\rm as}$, compared to the $42\pm 3~\mu{\rm as}$ value derived from the 2017 observations and published by the EHT Collaboration in 2019 Akiyama et al. (2019a, b, c, d). These unprecedented observations have, in turn, motivated extensive theoretical and numerical efforts to interpret the ring morphology and polarization signatures Psaltis et al. (2020); Cunha and Herdeiro (2020); Crinquand et al. (2022); Davelaar and Haiman (2022); Cunha et al. (2023); Galishnikova et al. (2023); Chen et al. (2025); Uniyal et al. (2025). In particular, the remarkable contributions from several groups Cunha and Herdeiro (2018); Mizuno et al. (2018); Gralla et al. (2019); Perlick and Tsupko (2022); Bronzwaer and Falcke (2021); Konoplya and Zhidenko (2021); Vagnozzi et al. (2023); Bambi et al. (2019); Chen et al. (2023); Ayzenberg et al. (2025), most notably the development of ray-tracing techniques Cunha et al. (2015, 2019); Hu et al. (2021); Bacchini et al. (2021); Zhang et al. (2025), have enabled the imaging of non-integrable and fully numerical black-hole spacetimes, providing a crucial bridge between theory and EHT observations.
@@
However, high-resolution backward ray-tracing simulations are notoriously computationally intensive, making it difficult to apply statistical inference methods such as Markov Chain Monte Carlo (MCMC) directly to constrain model parameters from observational data Johannsen et al. (2016). To overcome this limitation, we construct a general framework that unifies numerical imaging and parameter estimation in metric theories of gravity. Our approach employs an adaptive backward ray-tracing algorithm, inspired by the adaptive mesh refinement techniques in numerical relativity Clough et al. (2015); Buschmann et al. (2022), which reduces the computational scaling from geometric to nearly linear. This algorithm is combined with a Schwarzschild-calibrated normalization scheme that directly maps simulated image sizes to physical observables, independently of the observer’s configuration. Consequently, our framework enables high-precision shadow modeling at a minimal computational cost and, for the first time, allows statistically robust inference of black hole parameters from horizon-scale observations for arbitrary black hole spacetimes.
@@
In this Letter, we apply the above framework to the recently obtained Kerr–Bertotti–Robinson (Kerr–BR) black hole, an exact magnetized and rotating solution to the Einstein field equations Podolsky and Ovcharenko (2025). Unlike the Kerr-Melvin spacetimes Ernst and Wild (1976), the Kerr–BR spacetime belongs to the Petrov type-D class and features finite, asymptotically uniform electromagnetic fields and bounded ergoregions, thereby providing a more realistic model for astrophysical black holes embedded in external magnetic environments Ovcharenko and Podolský (2025); Wang (2025); Wang et al. (2025); Ali and Ghosh (2025); Zeng and Wang (2025); Zhang and Wei (2025). By matching the theoretical predictions of the Kerr–BR spacetime with the latest EHT observations of M87* and Sgr A*, we place quantitative constraints on the magnetic field strength. In particular, the analysis yields a magnetic field strength of $93.3^{+14.7}_{-23.8}\,\mathrm{G}$ for Sgr A*, consistent with the equipartition value $71\,\mathrm{G}$ inferred from polarized observations of submillimeter flares by the Atacama Large Millimeter/submillimeter Array (ALMA) Michail et al. (2023). Our results suggest that the magnetized environment can be consistently described by the exact Kerr–BR solution within general relativity (GR), without invoking any modification to the theory.
@@
with $I_{1}=1-B^{2}a^{2}/2$ and $I_{2}=1-B^{2}a^{2}$. The metric reduces to the Kerr solution in the limit $B=0$, but not to the Schwarzschild spacetime for $a=0$, reflecting the intrinsic coupling between rotation and the external magnetic field. Here $M$ is an integration constant interpreted as the mass parameter of the black hole. The outer and inner horizons of the black hole are determined by the roots of the metric function $\Delta$, namely $r_{\pm}\!=\!\tfrac{MI_{1}I_{2}\pm I_{1}\sqrt{M^{2}I_{2}-a^{2}I_{1}^{2}}}{I_{1}^{2}-B^{2}M^{2}I_{2}}$. The explicit form of the electromagnetic field is given in Ref. Podolsky and Ovcharenko (2025), but it is not essential to our present analysis; we are concerned only with its overall strength. In this Letter, our goal is to estimate the magnetic field strength of realistic supermassive black holes. To this end, we develop a new inference framework based on adaptive ray-tracing simulations. This framework directly links the observed shadow morphology to the strength of the underlying uniform magnetic field. Photon trajectories are integrated using the Hamiltonian formulation of null geodesics in the Kerr–BR spacetime,
@@
where $p_{\mu}=g_{\mu\nu}p^{\nu}$ and $p^{\mu}=(\dot{t},\dot{r},\dot{\vartheta},\dot{\varphi})$. The Killing vectors $\xi^{\mu}$ and $\psi^{\mu}$ yield two conserved quantities, the energy $\mathcal{E}$ and axial angular momentum $L_{z}$ Wald (1984), which allow us to numerically integrate the null geodesic equations in our backward ray-tracing scheme. Further computational details are given in Refs. Liu et al. (2025a, 2024a, b, 2024b, c). As illustrated in Fig. 1, a virtual observer is placed at a large radial distance, where the spacetime is effectively flat. The observer’s image plane, spanned by Cartesian coordinates $(x,y)$, is divided into an adaptive grid representing different initial photon directions within a field of view $\alpha_{\rm fov}$. Each grid point $(m,n)$ corresponds to a null geodesic that is integrated backward in time until the photon either crosses the event horizon or escapes to infinity. In our simulations, these two outcomes are distinguished on the image plane: escaping photons are marked by red pixels, while captured photons are marked by purple pixels. After all grid points have been traced, the resulting collection of purple pixels forms the simulated black hole shadow, whose outer boundary is illustrated by the purple ring in Fig. 1. For quantitative analysis, we introduce a reference circle, shown as the orange ring in Fig. 1, to compare with the noncircular shadow. This circle is defined as the circumcircle that passes through three characteristic points on the shadow boundary: $P_{A}(x_{A},y_{A})$ [^1] The normalized coordinates $(x_{i},y_{i})\in(0,1]$ are defined by the image resolution, e.g., $x_{A}=m/n$ for the $m$-th pixel in an $n\times n$ grid., $P_{B}(x_{B},y_{B})$, and $P_{C}(x_{C},y_{C})$. Meanwhile, a fourth point, $P_{D}(x_{D},y_{D})$, serves as a key indicator of the shadow’s deformation. Denoting the circumcenter coordinates by $P_{O}(x_{O},y_{O})$, the center satisfies the following system of equations:
@@
Through numerical backward ray-tracing simulations in the Kerr–BR spacetime, we establish explicit relationships linking observable quantities to black hole parameters and observer configuration. Accordingly, one can express $R_{s}\!\equiv\!R_{s}(\psi)$ and $\mathcal{D}_{s}\!\equiv\!\mathcal{D}_{s}(\psi)$ with $\psi\!\equiv\!\{\tilde{a},\tilde{B},\vartheta_{0}\}$, where $\tilde{a}\!=\!a/M$ and $\tilde{B}\!=\!BM$ denote the dimensionless spin and magnetic field strength, respectively, and $\vartheta_{0}$ is the observer’s inclination angle. For each parameter set $(\tilde{a},\tilde{B},\vartheta_{0})$, we ray-trace a two-dimensional shadow image on the observer’s image plane and extract the image-plane observables $(R_{s},\mathcal{D}_{s})$ from its pixels. Because $R_{s}$ is an image-plane quantity, we convert it into an apparent angular radius, $\theta_{s}$, by applying a Schwarzschild-based calibration such that $\theta_{s}\propto R_{s}$. Further details of this calibration are provided in Appendix A. With this calibration, the mappings $\theta_{s}(\psi)$ and $\mathcal{D}_{s}(\psi)$ provide a direct link between $\psi$ and the observable shadow features. This link allows us to use EHT measurements to constrain the black hole parameters $\tilde{a}$, $\tilde{B}$ and $\vartheta_{0}$. To achieve this, we construct the likelihood function as
@@
when both observables $\theta_{s}$ and $\mathcal{D}_{s}$ are available. Here, $D_{\mathrm{tot}}\!\equiv\!\{\theta_{\rm EHT},\sigma_{\theta},\mathcal{D}_{\rm EHT},\sigma_{\mathcal{D}}\}$ denotes a set of EHT observational data. In general, the probability densities $\mathcal{P}_{\theta_{s}}$ and $\mathcal{P}_{\mathcal{D}_{s}}$ are assumed to follow Gaussian distributions centered at the observed mean values $\theta_{\rm EHT}$ and $\mathcal{D}_{\rm EHT}$ with standard deviations $\sigma_{\theta}$ and $\sigma_{\mathcal{D}}$, respectively, when the reported uncertainties are symmetric. For asymmetric uncertainties, they are modeled as piecewise Gaussian distributions, with different standard deviations on each side of the mean. Then, the posterior distribution of theoretical parameters is obtained via Bayes’ theorem: $\mathcal{P}(\psi|D_{\mathrm{tot}})\propto\mathcal{L}(D_{\mathrm{tot}}|\psi)~\Pi(\psi)$, where $\Pi(\psi)$ represents the prior distribution, which we take to be uniform. We employ the MCMC method to sample from this posterior distribution and finally obtain the constraints on parameters $\tilde{a}$, $\tilde{B}$, and $\vartheta_{0}$.
Our overall modeling strategy is conceptually straightforward, but its implementation involves several key technical challenges. The first concerns how to map the pixel coordinates of the characteristic points $P_{A}$, $P_{B}$, $P_{C}$, and $P_{D}$ onto physically meaningful quantities relevant to astronomical observations, while the second relates to maintaining numerical precision and computational efficiency. The accuracy of the extracted shadow features is highly sensitive to the image resolution, as illustrated in Fig. 2 for the Kerr case where the axial ratio $\mathcal{D}_{s}$ varies with spin. The figure shows that numerical convergence is achieved only at sufficiently high resolutions and that the curves become progressively smoother. To address this, we develope an adaptive backward ray-tracing algorithm that performs high-resolution calculations only in the vicinity of the critical pixels. This adaptive scheme reduces the computational cost from exponential to nearly linear scaling with resolution, enabling us to explore the black hole parameter space at extremely high resolutions. The algorithmic details and validation tests of this method are presented in Appendix B.
@@
Results.—Having established a numerically consistent and efficient framework for shadow modeling, we next apply it to the EHT targets M87* and Sgr A*. This allows us to link their observed shadow features to theoretical predictions within the Kerr–BR spacetime and to quantitatively inferto quantitatively infer the underlying magnetic field strength $\tilde{B}$. To facilitate the comparison between observations and simulations, the relevant observational quantities and corresponding model parameters are listed in Table 1, with the latter expressed in geometrized units (c=G=M=1)(c\!\!=\!\!G\!\!=\!\!M\!\!=\!\!1).
@@
For the two EHT targets, M87* and Sgr A*, substituting the corresponding parameters from Table 1 into the above expression yields the theoretical angular radii $\theta_{s}$, which serve as the model predictions for subsequent parameter inference. Before comparing these theoretical predictions with observations, it is essential to ensure the numerical convergence of the simulated shadow contours. In general, the simulated shadow contours for different parameter choices converge at comparable resolutions, as illustrated in Fig. 2. Therefore, for calculating $\theta_{s}$ and $\mathcal{D}_{s}$, we find that an effective resolution of 4K (i.e., $4096\times 4096$ pixels) is sufficient to ensure full numerical convergence. With the numerical accuracy thus guaranteed, we proceed to constrain the model parameters using EHT observations. To ensure sufficient sampling of the parameter space, we generate $10^{4}$ simulated shadow images for each of M87* and Sgr A* under randomly selected combinations of $(\tilde{a},\tilde{B},\vartheta_{0})$. We then interpolate this dataset to obtain continuous mappings of $(\theta_{s},\mathcal{D}_{s})$, which are used to calculate the posterior distributions of theoretical parameters.
Figure 3: Corner plot showing the one- and two-dimensional posterior probability distributions of the BH parameters for M87*, based on the three-epoch combined measurements from the EHT Collaboration Akiyama et al. (2025). The dashed lines indicate the $68\%$ CL, while the dot-dashed line marks the mean value of $\tilde{B}$.
@@
In Fig. 3, we show the one- and two-dimensional posterior distributions of parameters for M87*, based on the three-epoch combined measurements of the EHT Collaboration Akiyama et al. (2025). The magnetic field strength is constrained to $BM\!=\!0.79^{+0.12}_{-0.11}\times 10^{-11}$. The posterior peaks at a high spin, a/M>0.759a/M\!>\!0.759, consistent with earlier jet-power arguments Akiyama et al. (2019d). The inclination angle remains weakly constrained, with $\vartheta_{0}\!>\!\frac{77\pi}{250}$ at the 68% confidence level (CL). Figure 4 presents the corresponding posterior distributions for Sgr A*, derived using the EHT 2022 observational data Akiyama et al. (2022). In this case, the term $\mathcal{P}_{\mathcal{D}_{s}}$ in Eq. (6) is omitted, since only $\theta_{s}$ is available (see Table 1). The inferred magnetic field strength is higher, $BM\!=\!1.53^{+0.22}_{-0.17}\times 10^{-11}$, while the correlations among $a/M$, BMBM, and $\vartheta_{0}$ appear marginal. This is not due to larger observational uncertainties but rather because the Sgr A* dataset provides no effective constraint on the shadow Axial ratio $\mathcal{D}_{s}$, leaving one fewer observable to break parameter degeneracies. Nevertheless, both analyses consistently favor the presence of a nonzero uniform magnetic field in the Kerr–BR spacetime, with the M87* results yielding a tighter credible interval for the field strength.
@@
$B_{\rm Gauss}=\frac{c^{4}}{G_{N}^{3/2}M}(BM)\approx 2.36\times 10^{19}\left(\frac{M_{\odot}}{M}\right)(BM)\ {\rm G},$ (7)
where $B_{\rm Gauss}$ is expressed in gauss. From the posterior distributions of $\tilde{B}$ and Eq. (7), we derive $B_{\rm Gauss}\!=\!0.0294^{+0.0045}_{-0.0055}\,{\rm G}$ (M87*) and $B_{\rm Gauss}\!=\!93.3^{+14.7}_{-23.8}\,{\rm G}$ (Sgr A*), both at 68% CL. Remarkably, the inferred field strength of Sgr A* agrees well with the equipartition magnetic field ($B_{\rm Gauss}\!\sim\!71{\rm G}$) derived from polarized ALMA observations of submillimeter flares Michail et al. (2023), while the much weaker field of M87* reflects the expected $1/M$ scaling of horizon-scale magnetization. This consistency provides compelling evidence that the Kerr-BR solution, an exact magnetized black hole spacetime in GR, can reproduce independent polarimetric magnetic field estimates when constrained purely by shadow geometry and EHT priors, even across systems differing in mass by over three orders of magnitude. It thus shows that horizo-scale magnetism can be self consistently explained within GR, without invoking any modification of Einstein’s theory, providing a unified geometric route to quantifying magnetic fields around both weakly and strongly magnetized black holes.
@@
Finally, we summarize all inferred constraints in Table 2. In addition to the results presented in Figs. 3 and 4, we incorporate other EHT measurements of M87* for comparison. These include the 2018 data reanalysis published in 2024 Akiyama et al. (2024) and the most recent three-epoch joint analysis that combines observations from 2017, 2018, and 2021 Akiyama et al. (2025). The latter study, which incorporated the 12-m Kitt Peak Telescope and the Northern Extended Millimetre Array, provided substantially improved baseline coverage and confirmed the long-term stability of the horizon-scale ring across all epochs. Taken together, these results demonstrate that the Kerr-BR model yields magnetic field strengths consistent with the latest EHT constraints while remaining fully compatible with GR, thereby reinforcing the viability of horizon-scale magnetism as a purely geometric observable.